OpenStudy (anonymous):

a/4+a/3=21 what is a?

OpenStudy (anonymous):

get lcd

OpenStudy (anonymous):

then add the fractions

OpenStudy (anonymous):

next, multiply both sides by lcd

OpenStudy (anonymous):

Lcd will be 12

OpenStudy (anonymous):

3a/12+4a/12=21 7a=21*12 a=3*12=36